OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (anonymous):

sin2x = -sinx hint: sin2x = 2sinxcosx

OpenStudy (anonymous):

Have you tried using the hint, then setting the two equations equal to one another?

OpenStudy (anonymous):

No I dont understand it at all can you please explain it?

OpenStudy (anonymous):

@hieuvo suggested using the identity sin (2x) = 2 sin x cos x So that would make your original equation 2 cos x sin x + sin x = 0, or 2 cos x sin x = -sin x

OpenStudy (anonymous):

Okay so how can I solve the equation?

OpenStudy (anonymous):

eliminate sinx

OpenStudy (kainui):

Are you sure the question isn't: sin^2(x)+sin(x)=0?

OpenStudy (anonymous):

That would change things a bit.

OpenStudy (kainui):

Yeah, and it is perfectly easily solvable too.

OpenStudy (anonymous):

the problem says sin2x + sin x = 0

OpenStudy (anonymous):

I mean sin^2x+ sin x =0

OpenStudy (kainui):

There are several ways you can do this. You can solve it like you would with the quadratic formula, or complete the square. You can substitute in something else for sinx so that it is more comfortable like the letter u. u^2+u=0 if you say u=sinx, then plug in sin(x) in the end to get your answers if that makes sense.

OpenStudy (anonymous):

Im sorry can you explain that a little more plaese?

OpenStudy (anonymous):

@Kainui suggested letting u = sin(x) Then your original problem becomes u^2 + u = 0

OpenStudy (kainui):

Just approach this like you would any old quadratic formula solving for x: ax^2+bx+c=0 except instead of x's you have sinx's that you're solving for.

OpenStudy (anonymous):

Oh okay i think I get it now

OpenStudy (anonymous):

Another way to think about this is to factor it, and use the zero product property.\[2 \cos x \sin x + \sin x = 0=\sin x(2\cos x+1) \implies \sin x =0~or~\cos x=-\frac{1}{2}\]This give us the solutions\[x \in \left\{ k \pi, \pm ( \frac{2 \pi}{3}+k \pi) \right\}~~ \forall k \in Z\]

OpenStudy (anonymous):

what is k?

OpenStudy (jennychan12):

sinx (sinx+1) = 0 sinx = 0 sinx = -1 solve for x

OpenStudy (anonymous):

Think of what angles, x, would cause sin(x) to be 0, or -1

OpenStudy (anonymous):

Would it be pi/2?

OpenStudy (anonymous):

pi/2 is 90 degrees. That is when sin(x) = 1. What angle causes sin(x) to be 0? And what angle causes sin(x) to be -1 ?

OpenStudy (jennychan12):

hint: unit circle :3

OpenStudy (anonymous):

180 degrees

OpenStudy (anonymous):

Right. sin(x) = 0 at 0 degrees and 180 degrees on the circle What about when sin(x) = 1?

OpenStudy (anonymous):

Remember \(\sin(x)\) is a periodic function with period \(2\pi\). This means if \(a\) is a solution, then \(\forall n \in \mathbb{N}\quad a+2\pi n\) is a solution.

OpenStudy (anonymous):

In degrees, the period is \(360^\circ \).

OpenStudy (anonymous):

\[ \forall n \in \mathbb{Z} \]

OpenStudy (anonymous):

I am confused now..can you please explain?

OpenStudy (anonymous):

First of all, did you get the normal solutions?

OpenStudy (anonymous):

were they 0, 180, and 90?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

My point is 0+360 is a solution too. 0+2*360 is also a solution. 0-360 is a solution. Does that make sense?

OpenStudy (anonymous):

okay so all of these are solutions?

OpenStudy (anonymous):

Yes, because if you rotate 360 degrees, you're back where you started.

OpenStudy (anonymous):

ohhhhhh okay

OpenStudy (anonymous):

So your solutions, plus any number of full rotations is a solution.

OpenStudy (anonymous):

So what you do is say that \(n\) is any integer... such as 0, 1, -1, 2, -2, etc.

OpenStudy (anonymous):

Then ALL solutions would be: \[ 0+360^\circ n, 90^\circ +360^\circ n, 180^\circ + 360^\circ n \]

OpenStudy (anonymous):

So originally you determined sin(x) = 0 sin(x) = -1 @wio represented all of the solutions for when sin(x) = 0. For what values does sin(x) = -1 ?

OpenStudy (anonymous):

It should have been sin(x) = -1, instead of sin(x) = 1 earlier

OpenStudy (anonymous):

oooh, whoops, looks like 90 isn't a solution.

OpenStudy (anonymous):

I think I mistyped it in an earlier post. Can't edit them without deleting it though.

OpenStudy (anonymous):

I really wish I could edit posts... I make little mistakes all the time.

OpenStudy (anonymous):

So the solutions would be 0, 180, 0 + 360?

OpenStudy (anonymous):

If we restrict our domain to one revolution around the circle, then x = 0 and x = 180 degrees are two solutions. The other solutions we would obtain by setting sin(x) = -1 For what value(s) of x does sin(x) = -1?

OpenStudy (anonymous):

is it 180?

OpenStudy (anonymous):

Nope... think of a circle